Rewrite the equation 2x – 8y = 48 as a function of x.
Answer by AnlytcPhil(1806)

Rewrite the equation 2x – 8y = 48 
as a function of x.

1. Solve for y
2. Replace y by f(x)

2x – 8y = 48
    -8y = 48 - 2x

Divide thru by -8 

      y = -6 + 1/4x

Replace y by f(x)

   f(x) = -6 + 1/4x

Edwin